# GeneMAN: Generalizable Single-Image 3D Human Reconstruction from Multi-Source Human Data

### v2

@@ 6 Conclusion
This appendix serves to further enrich the discourse established in our main paper. Sec. A offers additional information to supplement the related work. Sec. B presents the implementation details of GeneMAN. In Sec. C, we elaborate on the details of our multi-source human dataset. Sect. D provides supplementary experimental results, including more visualizations for the qualitative comparison, additional results of GeneMAN on in-the-wild images and CAPE [33], discussion on geometric quantitative evaluation, and ablation studies conducted on our multi-source human dataset. In Sec. E, we discuss the limitations of our approach.
Appendix A Supplement to Related Work
A.1 Text-to-3D Generation
The rapid advancements in 2D image generation have also significantly accelerated progress in the field of text-to-3D generation. DreamFusion [41] and Magic3D [25] have demonstrated that utilizing pretrained 2D text-to-image diffusion models [45, 44] as guidance can greatly enhance the optimization of 3D representations through score distillation sampling (SDS). Trained on large-scale 2D image datasets, these text-to-image models possess excellent capabilities in providing comprehensive 2D knowledge and hallucinating unseen scenes based on specific prompts. Nevertheless, as these models are limited to 2D knowledge, the aforementioned text-to-3D methods are susceptible to multi-view consistency issues. MVDream [53] improves 3D consistency by training a multi-view diffusion model that offers a robust multi-view prior. However, these methods are not well-suited for direct image-to-3D reconstruction tasks because images cannot be accurately captured through textual descriptions, leading to inconsistencies in color and texture across the generated assets.
A.2 Singe Image-to-3D Reconstruction
Single image-to-3D reconstruction has greatly benefited from the thriving advancement of diffusion models. Zero-1-to-3 [28] trains a view-dependent diffusion model on Objaverse [8] by explicitly incorporating the camera parameters into the diffusion process. It enables zero-shot image-conditioned novel view synthesis and facilitates general image-to-3D reconstruction with consistent geometric view. Magic123 [42] and DreamCraft3D [58] integrate the 3D prior from Zero-1-to-3 [28] with 2D prior from text-to-image diffusion model, employing a coarse-to-fine two-stage optimization to achieve high-quality 3D reconstruction. LRM [16] leverages both synthetic data from Objaverse [8] and multi-view data from MVImageNet [67] to train a transformer-based 3D architecture, enhancing its generalizability for 3D reconstruction. However, these general image-to-3D approaches often yield poor results in human reconstruction, resulting in inaccurate geometry of human bodies and the neglect of intricate details, such as facial features and clothing. This limitation arises from the lack of human-specific priors.
Appendix B Implementation Details
B.1 Camera Sampling
@@
For novel view guidance, we sample the camera distance from the range $\mathcal{U}(3.0,3.5)$. The elevation angle $\phi$ is drawn from $\mathcal{U}(-10^{\circ},45^{\circ})$, and the azimuth angle $\theta$ is uniformly sampled from $[-180^{\circ},180^{\circ})$. Additionally, the field of view (FOV) is constrained to the range $[20^{\circ},25^{\circ}]$, aligning with the fixed camera intrinsics optimized for our finetuned human diffusion models. Building upon HumanNorm [17], we segment the human body into four regions: the head, upper body, lower body, and full body. To enhance part-aware reconstruction with high-fidelity detail, we allocate a sampling probability of $0.7$ to the full body, and $0.1$ to each of the head, upper body, and lower body. We also manually adjust the camera distance for zooming in on specific parts. For instance, when refining the facial region, the camera distance is reduced to $\mathcal{U}(0.8,1.0)$. For half-body refinement (either upper or lower body), the camera distance is adjusted to $\mathcal{U}(1.5,2.0)$. Additionally, the camera center is shifted to ensure that the relevant keypoint aligns with the center of the rendered images. To achieve this, we use an off-the-shelf tool to identify 2D keypoints, which are then used during rasterization to back-project the coordinates into 3D space. Note that if a keypoint lies outside the image boundaries, we assign a sampling probability of zero to that point and renormalize the distribution accordingly.
B.2 Details of Each Stage
Geometry Initialization &\&& Sculpting.
@@ B.2 Details of Each Stage
In the phase of Geometry Initialization, we optimize Instant-NGP [37] from a resolution of $128$ to $384$ over the course of $5,000$ steps. The loss weights for this stage are set as follows: $\lambda_{\textnormal{r}}=1\times 10^{3},\lambda_{\textnormal{m}}=100,\lambda_{\textnormal{d}}=0.05,\lambda_{\textnormal{n}}=1,\lambda_{\textnormal{2D}}=0.1,\lambda_{\textnormal{3D}}=0.1$. Subsequently, we extract the resulting mesh from Instant-NGP with an isosurface resolution of $256$ as the geometry initialization. During the geometry sculpting stage, we refine the geometry adopting DMTet [52] at a resolution of $512$ for $3,000$ steps, enabling the capture of intricate details of humans. Inspired by HumanNorm [17], we incorporate a progressive positional encoding technique, where the mask on the position encoding for DMTet’s SDF features is gradually lifted to introduce higher-frequency components as training progresses. After $2,000$ iterations, the mask fully reveals all positions, allowing the encoding to capture both low- and high-frequency details. Empirically, we observe that omitting progressive positional encoding results in noisy surface reconstructions. The loss weights are set as follows: $\lambda_{r}=5\times 10^{3},\lambda_{vgg}=1\times 10^{3},\lambda_{sdf}=1.5\times 10^{3}$ and $\lambda_{sds}=1.0$. To ensure consistency between the rendered front view and the input image, we apply a relatively small guidance scale of $20$ for novel view generation using the normal- and depth-adapted diffusion models [17]. The noise timestep $t$ is sampled from $\mathcal{U}(0.02,0.8)$. Besides, we adopt the AdamW [31] optimizer with a base learning rate of $0.01$ during Geometry Initialization and $2\times 10^{-5}$ during Geometry Sculpting.
@@
Multi-Space Texture Refinement. For latent space texture refinement, we employ SDS optimization [41] in the latent space to optimize the coarse texture at an image resolution of $1024\times 1024$ for $10,000$ steps. The loss weights for the coarse texture stage are configured as follows: $\lambda^{color}_{rgb}=1\times 10^{3},\lambda^{color}_{m}=100,\lambda^{color}_{2D}=0.1,\lambda^{color}_{3D}=0.1$. Following this, we perform pixel-space texture refinement by optimizing the UV texture map for an additional 1,000 steps to achieve finer texture details. To ensure that the added noise does not corrupt the original content while retaining the capacity to enhance image details, we empirically set the starting timestep $t_{\textnormal{start}}$ to $0.05$ in Eq.8 in the main text. The weight for the VGG loss [54] in the pixel space texture refinement is set to $\lambda_{LP}=0.01$. We adopt the AdamW [31] optimizer with a base learning rate of $0.01$ in the coarse texture stage and $0.02$ in the fine texture stage.
Appendix C Details of Multi-Source Human Dataset
@@
RenderPeople Thuman2.0 Thuman3.0 HuMMan CostomHuman X-Humans Objaverse Human
1853 526 458 9072 647 3384 3388
@@
DNA Rendering ZJU-Mocap Neural Actor AIST++ Actor-HQ
8780 2646 4800 3853 3600
@@
2D Part Data DeepFashion LAION-5B
@@
To facilitate the training of multi-source human diffusion models, we construct a comprehensive multi-source human dataset containing $100K$ 2D human images and $52,345$ multi-view 3D human instances in total. Detailed statistics of our multi-source human dataset are provided in Tab. 2, with its composition described as follows: (1) The 2D human data consists of $20K$ human images from DeepFashion [29] and $80K$ human images filtered from LAION-5B [49]. The filtering process is conducted using YOLOv7 [60], eliminating images without human subjects or containing multiple individuals. Additionally, images are excluded if they have an aesthetics score below $4.5$, if the largest detected face measures less than $224\times 224$ pixels, or if the overall resolution falls below $640\times 1280$. To enhance semantic richness, each image is captioned using a finetuned BLIP model [23]. (2) The 3D human data is collected from 3D scanned human data, synthetic human data, video human data, and our synthetic human data. 3D Scanned Human Data contains human models sourced from the commercial dataset RenderPeople [1] and open-source datasets including CustomHumans [12], HuMMan [6], THuman2.0 [66], THuman3.0 [57] and X-Humans [51]. 3D Synthetic human data contains human-category objects filtered from Objaverse [8]. For both scanned and synthetic data, we adopt the dataset creation protocol of Zero-1-to-3 [28], with the exception that we uniformly select $48$ viewpoints across $360$ degrees in azimuth and set the resolution of the rendered image to $1024\times 1024$. Video Human Data contains five open-source datasets: DNA-Rendering [7], ZJU-MoCap [40], Neural Actor [27], AIST++ [22] and Actors-HQ [20]. For datasets containing background imagery, we apply CarveKit [50] to extract human silhouettes and produce corresponding RGBA images. When datasets feature multiple groups of surrounding cameras, we prioritize the group capturing the human subject at the image center. Note that we filter out cases of matting failures in these video datasets. As for DNA-Rendering [7], only Parts 1 and 2 of the released data are utilized. Our Synthetic Human Data contains human data synthesized by two augmentation techniques: ControlNet-based [68] image synthesis and image cropping synthesis. To further enhance the diversity of human identities and outfits in the dataset, inspired by En3D [34], we use ControlNet to synthesize a batch of multi-view human data based on diverse prompts describing outfits, genders, ages, and more, all generated by ChatGPT [2]. Unlike En3D, we extract templated or template-free poses and depth images from multi-view renderings of scanned human data (RenderPeople and XHumans [51]) as conditions fed into ControlNet and synthesize multi-view human images using diverse prompts. To ensure the quality of synthetic images, we concatenate pose and depth images horizontally across all views, generating multi-view humans in a single inference. This concatenation strategy ensures the texture consistency of the synthetic multi-view images. To handle the reconstruction of in-the-wild images with arbitrary human proportions, we crop the above fully body-length multi-view 3D human data into half-body length or three-quarters-body length images to create the augmented images. As for 3D data captioning, inspired by 3DTopia [15] and Cap3D [32], we use multi-modal large language model LLaVA [26] to generate captions for 3D objects by aggregating the descriptions from multiple views. The success of our GeneMAN framework demonstrates the effectiveness of our dataset in providing generalized priors for diverse human geometry and textures.
Appendix D Additional Experiment Results
D.1 More Qualitative Results
@@
For a more comprehensive geometric evaluation, we incorporated six additional SOTA human geometry reconstruction methods for comparison: PIFuHD [48], PaMIR [73], ICON [63], ECON [64], PHORHUM [4], and SIFU [72]. The geometric comparison results are presented in Fig. 17 and Fig. 18. It can be seen that our model effectively recovers natural human poses, accommodates loose clothing, and excels at reconstructing images with diverse body ratios. In addition, we provide more qualitative comparisons with state-of-the-art methods: PIFu [47], GTA [71], TeCH [18], SiTH [13] as shown in Fig. 19 to Fig. 25. Our method surpasses the compared methods in generating consistent, highly realistic textures with exceptional fidelity.
@@
Besides, in Fig. 13, we conduct additional comparisons with two generalizable methods: HumanLRM [61] and IDOL [74], both of which are feed-forward models trained on large-scale human datasets. As the code of Human-LRM is unavailable, we use the examples provided in its paper. Our model authentically reconstructs 3D humans that closely align with the front view, while IDOL fails to preserve facial details. Our model also produces natural poses with reasonable geometry (see the legs of the first and the third subjects in Fig. 13). Moreover, it achieves high-fidelity appearance with view-consistent details, significantly outperforming HumanLRM.
@@
We provide additional visualization results, with a particular focus on complex poses, in Fig. 15 (in-the-wild images) and Fig. 16 (CAPE [33]). These results highlight the strong generalization ability of GeneMAN to challenging poses.
@@
Figure 13: Additional Comparison on in-the-wild Cases with HumanLRM [61] and IDOL [74]. We conduct experiments on the examples reported by HumanLRM. Best view with zoomed in.
D.2 Quantitative Geometric Evaluation
@@
Following prior work, we perform a quantitative geometric comparison with baseline methods [47, 73, 63, 64, 71, 18, 13, 72] on the CAPE [33] dataset to assess geometry reconstruction quality. Specifically, we report two commonly used metrics: Chamfer distance (bi-directional point-to-surface) and Point-to-Surface (P2S) distance (1-directional point-to-surface), both measured in centimeters, between the ground truth scans and the reconstructed meshes. Additionally, to assess the fidelity of reconstructed local details, we calculate the $\mathcal{L}_{2}$ error between normal images rendered from the reconstructed and GT surfaces, referred to as Normal Consistency (NC). These renderings are obtained by rotating the camera around the surfaces at angles of $\{0^{\circ},120^{\circ},240^{\circ}\}$ relative to the frontal view.
@@
However, it is crucial to highlight that previous template-based methods [73, 63, 64, 71, 18, 13, 72] directly utilize ground truth SMPL [30, 39] provided in CAPE for evaluation. However, estimating body shape and pose parameters from a single image is an ill-posed problem due to ambiguity, leading to multiple possible solutions, as illustrated in Fig. 14. This presents an unfair advantage over our template-free approach. Moreover, our work focuses on reconstructing 3D humans with high fidelity from in-the-wild images, where accurate SMPL estimates are unavailable. To provide a fair comparison, we re-evaluate the baselines on the CAPE dataset under an inference mode, where GT body poses are not provided as input. The lack of access to such ground truth estimations leads to a notable performance drop for template-based methods, which deviates from the results presented in the original papers. The quantitative results are summarized in Tab. 3, where our method outperforms the compared approaches across all geometric metrics, demonstrating the superior reconstruction quality of our method.
@@
Figure 14: Visualization Results of the Reconstructed Geometry on the CAPE [33] dataset. While our template-free method successfully recovers plausible shapes, the orientation of the arms does not perfectly align with the ground truth, as highlighted by the red bounding box. This misalignment is difficult to avoid due to the inherent ambiguity, which can result in large 3D errors. Consequently, it is unfair to compare our approach with SMPL-based methods that utilize GT body parameters directly.
@@
Table 3: Geometric Comparison with State-of-the-art Methods on the CAPE [33] dataset. The best results are highlighted in bold. The second-place results are underlined.
[…]
PIFu 2.5580 2.5770 0.0925
PaMIR 2.5502 2.5920 0.0925
ICON 2.4147 2.4581 0.0872
ECON 2.0782 2.0296 0.0798
GTA 2.6785 2.7760 0.0914
TeCH 2.3217 2.4163 0.0935
SiTH 1.9182 2.0427 0.0726
SIFU 2.5226 2.5692 0.0875
@@
D.3 Ablation of Multi-Source Human Dataset
To validate the effectiveness of our constructed multi-source human dataset, we perform an ablation study on each component: 3D scan human data, video human data, augmented human data, and 2D human data. Specifically, we compare the reconstruction results using prior models trained with various combinations of these data components. The following experimental settings are evaluated: (a) “Baseline”, which replaces 2D and 3D prior models in GeneMAN with their original counterparts (Stable Diffusion V1.5 [44] and Zero-1-to-3 [28]); (b) “Baseline + 3D”, which uses prior models trained solely on 3D scanned human data; (c) “Baseline + 3D + Video”, which employs prior models trained on both 3D scanned human data and video human data; (d) “Baseline + 3D + Video + AUG” which incorporates prior models trained on 3D scan human data, video human data and augmented human data; (e) “Ours”, which utilizes prior models trained on the complete dataset. As detailed in Sec. 5.2 of the main text, we evaluate the performance of each setting using PSNR, LPIPS, and CLIP-Similarity across a set of $30$ test cases. As shown in Tab. 4, incorporating 3D scans significantly improves the model’s multi-view consistency, resulting in a 0.046 increase in CLIP-similarity. Furthermore, adding video data and augmented data enhances both texture quality and consistency. By utilizing the full dataset, our full-fledged method achieves the best performance in both texture quality and consistency.

### v3

@@ Acknowledgments
This appendix serves to further enrich the discourse established in our main paper. Sec. A presents the implementation details of GeneMAN. In Sec. B, we elaborate on the details of our multi-source human dataset. Sec. C provides supplementary experimental results, including more visualizations for the qualitative comparison, additional results of GeneMAN on in-the-wild images and CAPE Ma et al. (2020), additional quantitative comparisons across different body proportions, discussion on geometric quantitative evaluation, and ablation studies conducted on our multi-source human dataset. In Sec. D and E, we discuss the limitations of our approach, as well as the broader impacts and safeguards.
Appendix A Implementation Details
A.1 Camera Sampling
@@
For novel view guidance, we sample the camera distance from the range $\mathcal{U}(3.0,3.5)$. The elevation angle $\phi$ is drawn from $\mathcal{U}(-10^{\circ},45^{\circ})$, and the azimuth angle $\theta$ is uniformly sampled from $180^{\circ},180^{\circ})$. Additionally, the field of view (FOV) is constrained to the range $20^{\circ},25^{\circ}]$, aligning with the fixed camera intrinsics optimized for our finetuned human diffusion models. Building upon HumanNorm Huang et al. (2024a), we segment the human body into four regions: the head, upper body, lower body, and full body. To enhance part-aware reconstruction with high-fidelity detail, we allocate a sampling probability of $0.7$ to the full body, and $0.1$ to each of the head, upper body, and lower body. We also manually adjust the camera distance for zooming in on specific parts. For instance, when refining the facial region, the camera distance is reduced to $\mathcal{U}(0.8,1.0)$. For half-body refinement (either upper or lower body), the camera distance is adjusted to $\mathcal{U}(1.5,2.0)$. Additionally, the camera center is shifted to ensure that the relevant keypoint aligns with the center of the rendered images. To achieve this, we use an off-the-shelf tool to identify 2D keypoints, which are then used during rasterization to back-project the coordinates into 3D space. Note that if a keypoint lies outside the image boundaries, we assign a sampling probability of zero to that point and renormalize the distribution accordingly.
A.2 Details of Each Stage
Geometry Initialization &\& Sculpting.
@@ A.2 Details of Each Stage
In the phase of Geometry Initialization, we optimize Instant-NGP Müller et al. (2022) from a resolution of $128$ to $384$ over the course of $5,000$ steps. The loss weights for this stage are set as follows: $\lambda_{\textnormal{r}}=1\times 10^{3},\lambda_{\textnormal{m}}=100,\lambda_{\textnormal{d}}=0.05,\lambda_{\textnormal{n}}=1,\lambda_{\textnormal{2D}}=0.1,\lambda_{\textnormal{3D}}=0.1$. Subsequently, we extract the resulting mesh from Instant-NGP with an isosurface resolution of $256$ as the geometry initialization. During the geometry sculpting stage, we refine the geometry adopting DMTet Shen et al. (2021) at a resolution of $512$ for $3,000$ steps, enabling the capture of intricate details of humans. Inspired by HumanNorm Huang et al. (2024a), we incorporate a progressive positional encoding technique, where the mask on the position encoding for DMTet’s SDF features is gradually lifted to introduce higher-frequency components as training progresses. After $2,000$ iterations, the mask fully reveals all positions, allowing the encoding to capture both low- and high-frequency details. Empirically, we observe that omitting progressive positional encoding results in noisy surface reconstructions. The loss weights are set as follows: $\lambda_{r}=5\times 10^{3},\lambda_{vgg}=1\times 10^{3},\lambda_{sdf}=1.5\times 10^{3}$ and $\lambda_{sds}=1.0$. To ensure consistency between the rendered front view and the input image, we apply a relatively small guidance scale of $20$ for novel view generation using the normal- and depth-adapted diffusion models Huang et al. (2024a). The noise timestep $t$ is sampled from $\mathcal{U}(0.02,0.8)$. Besides, we adopt the AdamW Loshchilov (2017) optimizer with a base learning rate of $0.01$ during Geometry Initialization and $2\times 10^{-5}$ during Geometry Sculpting.
@@
Multi-Space Texture Refinement. For latent space texture refinement, we employ SDS optimization Poole et al. (2022) in the latent space to optimize the coarse texture at an image resolution of $1024\times 1024$ for $10,000$ steps. The loss weights for the coarse texture stage are configured as follows: $\lambda^{color}_{rgb}=1\times 10^{3},\lambda^{color}_{m}=100,\lambda^{color}_{2D}=0.1,\lambda^{color}_{3D}=0.1$. Following this, we perform pixel-space texture refinement by optimizing the UV texture map for an additional 1,000 steps to achieve finer texture details. To ensure that the added noise does not corrupt the original content while retaining the capacity to enhance image details, we empirically set the starting timestep $t_{\textnormal{start}}$ to $0.05$ in Eq.8 in the main text. The weight for the VGG loss Simonyan (2014) in the pixel space texture refinement is set to $\lambda_{LP}=0.01$. We adopt the AdamW Loshchilov (2017) optimizer with a base learning rate of $0.01$ in the coarse texture stage and $0.02$ in the fine texture stage.
Appendix B Details of Multi-Source Human Dataset
@@
RenderPeople 45 Thuman2.0 Yu et al. (2021) Thuman3.0 Su et al. (2022) HuMMan Cai et al. (2022)
1853 526 458 9072
CostomHuman Ho et al. (2023) X-Humans Shen et al. (2023) Objaverse Human Deitke et al. (2023)
647 3384 3388
@@
DNA Rendering Cheng et al. (2023) ZJU-Mocap Peng et al. (2021) Neural Actor Liu et al. (2021)
8780 2646 4800
AIST++ Li et al. (2021) Actors-HQ Işık et al. (2023)
3853 3600
@@
2D Part Data DeepFashion Liu et al. (2016) LAION-5B Schuhmann et al. (2022)
@@
To facilitate the training of multi-source human diffusion models, we construct a comprehensive multi-source human dataset containing $100K$ 2D human images and $52,345$ multi-view 3D human instances in total. Detailed statistics of our multi-source human dataset are provided in Tab. 2, with its composition described as follows: (1) The 2D human data consists of $20K$ human images from DeepFashion Liu et al. (2016) and $80K$ human images filtered from LAION-5B Schuhmann et al. (2022). The filtering process is conducted using YOLOv7 Wang et al. (2023), eliminating images without human subjects or containing multiple individuals. Additionally, images are excluded if they have an aesthetics score below $4.5$, if the largest detected face measures less than $224\times 224$ pixels, or if the overall resolution falls below $640\times 1280$. To enhance semantic richness, each image is captioned using a finetuned BLIP model Li et al. (2024). (2) The 3D human data is collected from 3D scanned human data, synthetic human data, video human data, and our synthetic human data. 3D Scanned Human Data contains human models sourced from the commercial dataset RenderPeople 45 and open-source datasets including CustomHumans Ho et al. (2023), HuMMan Cai et al. (2022), THuman2.0 Yu et al. (2021), THuman3.0 Su et al. (2022) and X-Humans Shen et al. (2023). 3D Synthetic human data contains human-category objects filtered from Objaverse Deitke et al. (2023). For both scanned and synthetic data, we adopt the dataset creation protocol of Zero-1-to-3 Liu et al. (2023), with the exception that we uniformly select $48$ viewpoints across $360$ degrees in azimuth and set the resolution of the rendered image to $1024\times 1024$. Video Human Data contains five open-source datasets: DNA-Rendering Cheng et al. (2023), ZJU-MoCap Peng et al. (2021), Neural Actor Liu et al. (2021), AIST++ Li et al. (2021) and Actors-HQ Işık et al. (2023). For datasets containing background imagery, we apply CarveKit Selin (2023) to extract human silhouettes and produce corresponding RGBA images. When datasets feature multiple groups of surrounding cameras, we prioritize the group capturing the human subject at the image center. Note that we filter out cases of matting failures in these video datasets. As for DNA-Rendering Cheng et al. (2023), only Parts 1 and 2 of the released data are utilized. Our Synthetic Human Data contains human data synthesized by two augmentation techniques: ControlNet-based Zhang et al. (2023) image synthesis and image cropping synthesis. To further enhance the diversity of human identities and outfits in the dataset, inspired by En3D Men et al. (2024), we use ControlNet to synthesize a batch of multi-view human data based on diverse prompts describing outfits, genders, ages, and more, all generated by ChatGPT Achiam et al. (2023). Unlike En3D, we extract templated or template-free poses and depth images from multi-view renderings of scanned human data (RenderPeople and XHumans Shen et al. (2023)) as conditions fed into ControlNet and synthesize multi-view human images using diverse prompts. As shown in Fig. 11 and Fig. 12, examples of Image Cropping Synthesis and ControlNet-based Synthesis are presented, respectively. To ensure the quality of synthetic images, we concatenate pose and depth images horizontally across all views, generating multi-view humans in a single inference. This concatenation strategy ensures the texture consistency of the synthetic multi-view images. To handle the reconstruction of in-the-wild images with arbitrary human proportions, we crop the above fully body-length multi-view 3D human data into half-body length or three-quarters-body length images to create the augmented images. As for 3D data captioning, inspired by 3DTopia Hong et al. (2024) and Cap3D Luo et al. (2024), we use multi-modal large language model LLaVA Liu et al. (2024) to generate captions for 3D objects by aggregating the descriptions from multiple views. The success of our GeneMAN framework demonstrates the effectiveness of our dataset in providing generalized priors for diverse human geometry and textures.
Appendix C Additional Experiment Results
C.1 More Qualitative Results
@@
For a more comprehensive geometric evaluation, we incorporated six additional SOTA human geometry reconstruction methods for comparison: PIFuHD Saito et al. (2020), PaMIR Zheng et al. (2021), ICON Xiu et al. (2022), ECON Xiu et al. (2023), PHORHUM Alldieck et al. (2022), and SIFU Zhang et al. (2024c). The geometric comparison results are presented in Fig. 17 and Fig. 18. It can be seen that our model effectively recovers natural human poses, accommodates loose clothing, and excels at reconstructing images with diverse body ratios. In addition, we provide more qualitative comparisons with state-of-the-art methods: PIFu Saito et al. (2019), GTA Zhang et al. (2024b), TeCH Huang et al. (2024b), SiTH Ho et al. (2024) as shown in Fig. 19 to Fig. 25. Our method surpasses the compared methods in generating consistent, highly realistic textures with exceptional fidelity.
@@
Besides, in Fig. 13, we conduct additional comparisons with two generalizable methods: HumanLRM Weng et al. (2024) and IDOL Zhuang et al. (2024), both of which are feed-forward models trained on large-scale human datasets. As the code of Human-LRM is unavailable, we use the examples provided in its paper. Our model authentically reconstructs 3D humans that closely align with the front view, while IDOL fails to preserve facial details. Our model also produces natural poses with reasonable geometry (see the legs of the first and the third subjects in Fig. 13). Moreover, it achieves high-fidelity appearance with view-consistent details, significantly outperforming HumanLRM.
@@
We provide additional visualization results, with a particular focus on complex poses, in Fig. 15 (in-the-wild images) and Fig. 16 (CAPE Ma et al. (2020)). These results highlight the strong generalization ability of GeneMAN to challenging poses.
@@
Figure 13: Additional Comparison on in-the-wild Cases with HumanLRM Weng et al. (2024) and IDOL Zhuang et al. (2024). We conduct experiments on the examples reported by HumanLRM. Best view with zoomed in.
C.2 Additional Quantitative Comparison Different Across Body Proportions
To assess the robustness of our method under varying human proportions, we additionally evaluate it on in-the-wild images that include head-shot, half-body, and full-body inputs. For comparison, we only consider the template-free method PIFu, since template-based approaches generally fail to reconstruct head-shot and half-body inputs reliably. As shown in Table 3, GeneMAN consistently outperforms PIFu across all proportions setting, demonstrating its strong generalization capability.
Table 3: Quantitative Comparison Across Different Human Proportions. The best results are highlighted in bold.
Method Full-body Half-body Head-shot
PSNR $\uparrow$ LPIPS $\downarrow$ CLIP-Sim $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ CLIP-Sim $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ CLIP-Sim $\uparrow$
PIFu 28.039 0.029 0.622 25.383 0.052 0.477 25.047 0.075 0.522
GeneMAN (Ours) 32.529 0.012 0.751 28.245 0.016 0.681 26.374 0.022 0.661
C.3 Quantitative Geometric Evaluation
@@
Following prior work, we perform a quantitative geometric comparison with baseline methods Saito et al. (2019); Zheng et al. (2021); Xiu et al. (2022, 2023); Zhang et al. (2024b); Huang et al. (2024b); Ho et al. (2024); Zhang et al. (2024c) on the CAPE Ma et al. (2020) dataset to assess geometry reconstruction quality. Specifically, we report two commonly used metrics: Chamfer Distance (CD) and Point-to-Surface distance (P2S), both measured in centimeters, between the ground truth scans and the reconstructed meshes. Additionally, to assess the fidelity of reconstructed local details, we calculate the $\mathcal{L}_{2}$ error between normal images rendered from the reconstructed and GT surfaces, referred to as Normal Consistency (NC). These renderings are obtained by rotating the camera around the surfaces at angles of $\{0^{\circ},120^{\circ},240^{\circ}\}$ relative to the frontal view.
@@
However, it is crucial to highlight that previous template-based methods Zheng et al. (2021); Xiu et al. (2022, 2023); Zhang et al. (2024b); Huang et al. (2024b); Ho et al. (2024); Zhang et al. (2024c) directly utilize ground truth SMPL Loper et al. (2023); Pavlakos et al. (2019b) provided in CAPE for evaluation. However, estimating body shape and pose parameters from a single image is an ill-posed problem due to ambiguity, leading to multiple possible solutions, as illustrated in Fig. 14. This presents an unfair advantage over our template-free approach. Moreover, our work focuses on reconstructing 3D humans with high fidelity from in-the-wild images, where accurate SMPL estimates are unavailable. To provide a fair comparison, we re-evaluate the baselines on the CAPE dataset under an inference mode, where GT body poses are not provided as input. The lack of access to such ground truth estimations leads to a notable performance drop for template-based methods, which deviates from the results presented in the original papers. The quantitative results are summarized in Tab. 4, where our method outperforms the compared approaches across all geometric metrics, demonstrating the superior reconstruction quality of our method.
@@
Figure 14: Visualization Results of the Reconstructed Geometry on the CAPE Ma et al. (2020) dataset. While our template-free method successfully recovers plausible shapes, the orientation of the arms does not perfectly align with the ground truth, as highlighted by the red bounding box. This misalignment is difficult to avoid due to the inherent ambiguity, which can result in large 3D errors. Consequently, it is unfair to compare our approach with SMPL-based methods that utilize GT body parameters directly.
@@
Table 4: Geometric Comparison with State-of-the-art Methods on the CAPE Ma et al. (2020) dataset. The best results are highlighted in bold. The second-place results are underlined.
Methods CD $\downarrow$ P2S $\downarrow$ NC $\downarrow$
PIFu Saito et al. (2019) 2.5580 2.5770 0.0925
PaMIR Zheng et al. (2021) 2.5502 2.5920 0.0925
ICON Xiu et al. (2022) 2.4147 2.4581 0.0872
ECON Xiu et al. (2023) 2.0782 2.0296 0.0798
GTA Zhang et al. (2024b) 2.6785 2.7760 0.0914
TeCH Huang et al. (2024b) 2.3217 2.4163 0.0935
SiTH Ho et al. (2024) 1.9182 2.0427 0.0726
SIFU Zhang et al. (2024c) 2.5226 2.5692 0.0875
@@
C.4 Ablation of Multi-Source Human Dataset
To validate the effectiveness of our constructed multi-source human dataset, we perform an ablation study on each component: 3D scan human data, video human data, augmented human data, and 2D human data. Specifically, we compare the reconstruction results using prior models trained with various combinations of these data components. The following experimental settings are evaluated: (a) “Baseline", which replaces 2D and 3D prior models in GeneMAN with their original counterparts (Stable Diffusion V1.5 Rombach et al. (2022) and Zero-1-to-3 Liu et al. (2023)); (b) “Baseline + 3D”, which uses prior models trained solely on 3D scanned human data; (c) “Baseline + 3D + Video", which employs prior models trained on both 3D scanned human data and video human data; (d) “Baseline + 3D + Video + AUG" which incorporates prior models trained on 3D scan human data, video human data and augmented human data; (e) “Ours”, which utilizes prior models trained on the complete dataset. As detailed in Sec. 5.2 of the main text, we evaluate the performance of each setting using PSNR, LPIPS, and CLIP-Similarity across a set of $30$ test cases. As shown in Tab. 5, incorporating 3D scans significantly improves the model’s multi-view consistency, resulting in a 0.046 increase in CLIP-similarity. Furthermore, adding video data and augmented data enhances both texture quality and consistency. By utilizing the full dataset, our full-fledged method achieves the best performance in both texture quality and consistency.